answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

What is the probability of getting heads more than once if you flip a coin three times?

For 3 coin flips:

87% chance of getting heads at least once

25% chance of getting heads twice

13% chance of getting heads all three times

What is the probability that at least 2 students in a class of 36 have the same birthday?

About 83.2%

The probability that non of the 36 students have the same birthday (not considering

February 28 of the leap year) is given by the following relation:

P(non out of n have same bd) = Π1n-1 [(365-i)/365]

P(non out of 36 have same bd) = (364/365)(363/365)(362/365) ... (331/365)(330/365) =

= 0.167817892.. ≈ 16.8%

So the probability of at least 2 having the same birthday is about 1 - .168 = 0.832 =

83.2%

What is the probability drawing a black king from a pack of cards?

Assuming there are no Joker cards the chance is one in twenty six.

There are fifty two cards in a pack and only two of them are black kings.

How do you solve if you throw a 6 sided die 84 times how many time would you expect 4 to come up?

If it is a fair die, the probability of throwing a 4 in a single go is 1/6. That is, you expect to get a 4 in 1/6 of the throws. So for 84 throws you would expect to get a 4 in 84*(1/6) = 14 times.

What is the probability of rolling two even numbers on two number cubes?

1/2 or 50%

Answer:

If the question means rolling even numbers on each at the same time (which it appears to) then the previous answer is incorrect.

There's a 50% chance on each die of getting an even number, but since we require them to occur together the individual probabilities are multiplied, giving 25%.

This is easy to illustrate:

Die 1 Die 2

Odd Odd

Odd Even

Even Odd

Even Even

In only 1 of the four possibilities does evens come up on both.

What sample size is sufficient for stat?

A sample size of one is sufficient to enable you to calculate a statistic.

The sample size required for a "good" statistical estimate will depend on the variability of the characteristic being studied as well as the accuracy required in the result. A rare characteristic will require a large sample. A high degree of accuracy will also require a large sample.

What is the probability that you toss a head or a tail?

Assuming a two-sided coin, and that you make the the toss, the probability of tossing a head or a tail is 100%. The probability of tossing a head is 50%. The probability of tossing a tail is 50%.

What is the probaility of drawing a 10 from a standard deck?

number of cards that are 10 =4

number of cards in a standard deck =52

Probability of drawing a 10 =4/52 = 1/13

When to you use a z scores or t scores?

If the distribution is Gaussian (or Normal) use z-scores. If it is Student's t, then use t-scores.

What sum are you least likely to get rolling two dice?

all numbers on a die have the same probability to land so any number between 2 and 12 are just as likely as any other number between 2 and 12, just as any other die hase the same probability 2 d20, 2-40 any to are just as likely as the rest. so 0 is the least likely number.

Well, yes, since 0 is impossible to be the sum of 2 dice, it is least likely. So is any other number less than 2 or greater than 12, or a number that is not a whole number. But of those that actually can come up as the sum of 2 dice, 7 is the most likely because there are 6 ways to get it (1 and 6, 2 and 5, 3 and 4, 4 and 3, etc.). The least likely sums are 2 and 12, because each can only occur 1 way (a pair of 1's or a pair of 6's). These are referred 2 by gamblers as snake eyes and box cars.

What is the probability of getting 2 heads when you toss three coins?

If you mean 'at least' 2 heads, the probability is 50%. If you mean exactly 2,

the probability is 3/8, or 37.5%.

There are 3 independent coin tosses, each of which is equally likely to come up

heads or tails. That's a total of 2 * 2 * 2 or 8 possible outcomes (HHH, HHT, HTH,

etc.). Of these, 4 include 2 or 3 heads, which is half of 8. Only 3 include exactly 2

heads, so the probability of that is 3/8.